ND-GAr: HPgTPC+ECAL Weekly Meeting

Disappearance Room (WH13SE) and Zoom

Disappearance Room (WH13SE) and Zoom


DUNE Phase 2 meetings (ND an FD) starting biweekly on Mondays.  We would like to have regular updates on various test stand efforts.  Let us know if you want to give an update on light collection, GOAT, TOAD, etc.


There are minutes attached to this event. Show them.
    • 10:00 AM 10:20 AM
      GArsoft Update 20m
      Speaker: Leo Bellantoni (FNAL)
    • 10:25 AM 10:45 AM
      Roundtable Updates 20m
      Speakers: Alfons Weber (JG-Uni Mainz & Fermilab), Diego Gonzalez Diaz, Naseem Khan, Tanaz Mohayai (Fermilab)


      Tanaz:  Trying to wrap up IROC data collection before moving to GEMs.  Taking some missing data with Ar/CH4 and maybe Ar/CO2.  New problems with CREMAT card, maybe related to calibration pulses.  But not limiting data collection right now.

      Alan:  First noticed it last Friday.  Big pulses cause electronics to oscillate.  Can see this when with HV off.  Seems coming from new 672Ortec shaper.  Did not see it with old module and it was working fine earlier with this shaper.  Can switch back to older shaper, but noisier.  But probably low enough to switch back to old module.  Would also like to see if it can operate in for a week.

      Leo:  Asks about moving GOAT to Helen Edwards Building.  Tanaz: Wants to finish GEM tests before moving the vessel.  Move takes a while.  But next year might move there.



      Tanaz: Clean room space secured at SciDet for bench test at 1 atm.  Want to speed up data collection. Need to figure out how to secure a gas bottle.  GEM has a source holder. 

      Diego: Asks about resistor divider.  Tanaz:  At the moment yes, and each GEM layer gets the same voltage.  Diego suggests talking to the CERN people at some point and is also happy to give some feedback.  Patrick also asks about pixelating, and can we couple them to the SAMPA electronics in at some point.  Right now all of the GEM pads are read out and there is only an ethernet cable (from the company).



      Naseem: ORC successful on last Thursday. Can run to voltage and fill to pressure.  They just want to have a look after electronics are in.  Focusing on putting electronics in this week and understanding the mapping.

      Patrick:  Electronics are all tested. 


      Light Collection

      Diego:  Had done a long-exposure measurement of optical gain.  Measurement with single thick GEM at 1 bar is 1000-2000.  But it developed a high current during operation.  They are cleaning it and dark current going down.   Developing a frame for glass GEMs from Liverpool.  Will proceed with double thick GEM readout.   Also plan to order ne acrylic thick GEMs from CERN.   Also simulation progress and can simulate tracks. 



      Tom:  Our art version is about a year old.  Takes a week to update by hand, but they might have scripts that can make it go faster.